OpenStudy (anonymous):

If f(x)=sqrtx^2-1 and g(x) = sqrt x-1, which expression represent f(x)/g(x) for x >1

OpenStudy (anonymous):

\[f(x) =\sqrt{x^2-1} and g(x)=\sqrt{x-1}, which expression represent \frac{f(x) }{ g(x) for x }\]

OpenStudy (anonymous):

for x>1